OpenStudy (anonymous):

solve the following system by substitution: 3x + y - 2z = 22 x + 5y + z = 4 x = -3z I've tried a few ways to solve, but I can't get it. Can anyone show me a way to solve?

mathslover (mathslover):

Let us use the third equation first. We have : \(\color{blue}{x = -3z}\) according to the third equation. Right?

mathslover (mathslover):

Now, we will plug-in the value of x from third equation, in any of the two equations. Let us pick the second one as it will be simpler for us to go through it. We have : \(x + 5y +z = 4\) , according to the second equation. \(\color{blue}{-3z} + 5y + z = 4\) Note that I plugged in the value of x from third equation into the second equation/ Can you simplify the above equation a bit now?

OpenStudy (anonymous):

Okay, I solved for z but I just got \[y=\frac{ 2(z+2) }{ 5 }\]

mathslover (mathslover):

Okay, so, you solved for y. Cool! So, put this value of y : \( \color{red}{y = \cfrac{2(z+2)}{5} }\) and that of x : \(\color{blue}{x = -3z}\) in equation (1). See what you get.

OpenStudy (anonymous):

oops I meant I solved for y not z yes haha okay, so I'm solving this for z: 3(-3z) + 2(z+2)/5 - 2z = 22 and I got z = -2. Am I on the right track?

mathslover (mathslover):

Excellent! You're on the right track. Now plug-in this value of z in : \(\color{blue}{x = -3z }\) and thus, you will get the value of x.

OpenStudy (anonymous):

So x=6!

OpenStudy (anonymous):

And now I can plug my x and z values back into the first equation to solve for my y value?

OpenStudy (anonymous):

okay, from 3(6)+y-2(-2)=22 I got y=0

mathslover (mathslover):

Just plug-in the value of z in : \(\color{red}{y = \cfrac{2(z+2)}{5}}\) Yeah, that's right. Great work.

OpenStudy (anonymous):

Awesome! Thank you so much for everything, you made that seem way easier than I thought it was going to be!!

mathslover (mathslover):

:) You're welcome @boldlygo ! Though, I just gave you a hint and you did the rest. So, applauds for you. Keep it up and welcome to OpenStudy!